# Standard Lyndon loop words: weighted orders

### v1

@@ 4.3. Affine Weyl groups
(4.10) $s_{i}(\mu)=\mu-(\alpha_{i},\mu)\alpha_{i}^{\vee}\ \qquad\text{and}\qquad s_{i}(\lambda)=\lambda-(\lambda,\alpha^{\vee}_{i})\alpha_{i}\,.$
@@
(4.12) $s_{i}(\lambda,d)=\left(\lambda-(\lambda,\alpha^{\vee}_{i})\alpha_{i},d\right)\quad\forall\,i\in I\,,$
(4.13) $s_{0}(\lambda,d)=\left(\lambda-(\lambda,\theta^{\vee})\theta,d+(\lambda,\theta^{\vee})\right)$
@@ 4.4. Reduced decompositions
Let ${\mathbb{N}}={\mathbb{Z}}_{\geq 0}$. Recall that the length of an element $x\in\widehat{W}$, denoted by $l(x)\in{\mathbb{N}}$, is the smallest number $l\in{\mathbb{N}}$ such that we can write $x=s_{i_{1-l}}\dots s_{i_{0}}$ for various $i_{1-l},\dots,i_{0}\in\widehat{I}$. Every such factorization is called a reduced decomposition of $x$. Given such a reduced decomposition, the terminal subset of the affine root system is:
@@ 4.6. Identification of two orders
give convex orders of the sets $\Delta^{+}\times{\mathbb{Z}}_{<0}$ and $\Delta^{+}\times{\mathbb{Z}}_{\geq 0}$, respectively. We note that if $\beta_{k}=(\alpha,d)$ and $\beta_{k+l}=(\alpha^{\prime},d^{\prime})$, then
(4.27) $\beta_{k+l}=\widehat{\mu}(\beta_{k})\qquad\Longrightarrow\qquad\alpha=\alpha^{\prime}\text{ and }d=d^{\prime}+(\alpha,\mu)\,,$
due to (4.17). This reveals a periodicity of the entire set $\Delta^{+}\times{\mathbb{Z}}$.
@@
For any $(\alpha,d)\in L_{N}$, the relative exponent of the first letter in $\ell(\alpha,-d)$ lies in $(-N-1;N]$, due to Corollary 3.11 and Proposition 3.15. Thus, for any $(\alpha,d)\in L_{M}$, $(\beta,e)\in L_{N}$ with $M>N$, we have $\ell(\alpha,-d)>\ell(\beta,-e)$. As for the affine roots from the same block, consider $\beta_{r-Nl},\beta_{s-Nl}\in L_{N}$ with $1-l\leq s<r\leq 0$. If $\beta_{r}=(\alpha,d)$ and $\beta_{s}=(\beta,e)$, then $\beta_{r-Nl}=(\alpha,d+N\cdot f(\alpha))$ and $\beta_{s-Nl}=(\beta,e+N\cdot f(\beta))$, due to (4.27). On the other hand, the words $\ell(\alpha,-d-N\cdot f(\alpha))$ and $\ell(\beta,-e-N\cdot f(\beta))$ are obtained from $\ell(\alpha,-d)$ and $\ell(\beta,-e)$, respectively, by decreasing each relative exponent by $N$, due to Proposition 3.15. Since the latter operation obviously preserves the lexicographic order, and $\ell(\alpha,-d)<\ell(\beta,-e)$ as a consequence of $r>s$, we obtain the required inequality $\ell(\alpha,-d-N\cdot f(\alpha))<\ell(\beta,-e-N\cdot f(\beta))$. ∎
@@
(4.31) $\dots<\beta_{3}<\beta_{2}<\beta_{1}<\beta_{0}<\beta_{-1}<\beta_{-2}<\cdots$
@@
(4.32) $L_{<(i,d)}=\Big{\{}(\alpha,p)\,\Big{|}\,\alpha\in\Delta^{+},\,p\in{\mathbb{Z}}_{\geq 0},\,\ell(\alpha,-p)<\ell(\alpha_{i},-d)\Big{\}}.$
@@
First, let us prove that $\ell(\alpha,-p)<\ell(\alpha_{i},-d)=i^{(-d)}$ implies $p<p(\alpha)$. Let $j^{(-e)}$ be the first letter of $\ell(\alpha,-p)$, so that $j^{(-e)}<i^{(-d)}$. Hence, $e/c_{j}\leq d/c_{i}$ and the inequality is strict if $j\geq i$. This is equivalent to $e<p_{j}$, due to the definition (4.33). Then for any letter $\imath^{(-s)}\in\ell(\alpha,-p)$, we have $\imath^{(-s+1)}\leq j^{(-e)}\leq i^{(-d)}$ with the first inequality due to Theorem 3.6. As above, this implies $s-1<p_{\imath}$, so that $s\leq p_{\imath}$. Summing all these inequalities, we obtain the desired inequality $p<p(\alpha)$.
Let us prove the opposite implication by contradiction: assume that $\ell(\alpha,-p)>\ell(\alpha_{i},-d)$ for some $\alpha\in\Delta^{+}$ and $p<p(\alpha)$. Let $j^{(-e)}$ be the first letter of $\ell(\alpha,-p)$, so that $j^{(-e)}\geq i^{(-d)}$. We consider two cases:
  1. $\bullet$
    Case 1: $j^{(-e)}=i^{(-d)}$.
    As $\ell(\alpha,-p)$ is Lyndon, any letter $\imath^{(-s)}\in\ell(\alpha,-p)$ satisfies $j^{(-e)}\leq\imath^{(-s)}$. Therefore, $s/c_{\imath}\geq e/c_{j}=d/c_{i}$ and the inequality is strict for $\imath<i$. Thus, $s\geq p_{\imath}$. Summing all these inequalities, we obtain $p\geq p(\alpha)$, a contradiction.
  2. $\bullet$
    Case 2: $j^{(-e)}>i^{(-d)}$.
    As in Case 1, any letter $\imath^{(-s)}\in\ell(\alpha,-p)$ satisfies $\imath^{(-s)}\geq j^{(-e)}>i^{(-d)}$. Hence, $s/c_{\imath}\geq d/c_{i}$ and the inequality is strict for $\imath<i$. Thus, $s\geq p_{\imath}$. Summing all these inequalities, we again obtain $p\geq p(\alpha)$, a contradiction.
This completes our proof of $\ell(\alpha,-p)<\ell(\alpha_{i},-d)$ for any $0\leq p<p(\alpha)$. ∎
@@ 5.1. Quantum groups
If we write $\varphi_{i}=q_{i}^{h_{i}}$ and take the limit $q\rightarrow 1$, then $U_{q}({\mathfrak{g}})$ degenerates to $U({\mathfrak{g}})$. It is thus natural that many features of the latter also admit $q$-deformations. For example, let us recall the notion of standard Lyndon words from Subsections 2.12.7, and consider the following $q$-version of the construction of (2.4) and Definition 2.8.
@@
(5.6) $U_{q}({\mathfrak{n}}^{+})\ =\bigoplus^{k\in{\mathbb{N}}}_{\ell_{1}\geq\dots\geq\ell_{k}\text{ standard Lyndon words}}{\mathbb{Q}}(q)\cdot e_{\ell_{1}}\dots e_{\ell_{k}}\ =\\ \bigoplus_{w\text{--standard words}}{\mathbb{Q}}(q)\cdot e_{w}\,.$
@@ 5.5. Quantum loop algebras
(5.12) $\text{hdeg }x=\alpha\qquad\text{and}\qquad\text{vdeg }x=d\,,$
@@
(5.16) $e_{w}=e_{\ell_{1}}\dots e_{\ell_{k}}\qquad\text{and}\qquad f_{w}=f_{\ell_{1}}\dots f_{\ell_{k}}$
@@
The analogous result also holds with $L{\mathfrak{n}}^{+}\leftrightarrow L{\mathfrak{n}}^{-}$ and $e\leftrightarrow f$.
@@ 5.9. Quantum affine algebras
We may define the root vectors:
@@
(5.26) $\Omega\colon e_{i}\mapsto f_{i},\ f_{i}\mapsto e_{i},\ \varphi^{\pm 1}_{i}\mapsto\varphi_{i}^{\mp 1},\ q\mapsto q^{-1}\qquad\forall\,i\in\widehat{I}\,.$
@@
(5.29) $\not\exists\ \tilde{\alpha}^{\prime},\tilde{\beta}^{\prime}\in\widehat{\Delta}^{\mathrm{re},+}\quad\text{s.t.}\quad\tilde{\alpha}<\tilde{\alpha}^{\prime}<\tilde{\beta}^{\prime}<\tilde{\beta}\quad\text{and}\quad\tilde{\alpha}+\tilde{\beta}=\tilde{\alpha}^{\prime}+\tilde{\beta}^{\prime}\,.$
@@ 5.13. Interplay of two algebras
It suffices to prove only (5.35) since (5.36) would then follow as $\Omega$ commutes with the extended affine braid group action (due to (5.27) and $\Omega\circ T_{\tau}=T_{\tau}\circ\Omega$ for $\tau\in{\mathcal{T}}$).
@@
$T_{\widehat{\omega_{i,d}}}=T^{p_{i}}_{\widehat{\omega_{i}^{\vee}}}\cdot\prod_{j\neq i}T^{p_{j}}_{\widehat{\omega_{j}^{\vee}}}=T^{d}_{\widehat{\omega_{i}^{\vee}}}\cdot\prod_{j\neq i}T^{p_{j}}_{\widehat{\omega_{j}^{\vee}}}\,.$
@@ 5.15. PBW-type bases via quarter subalgebras
(5.39) $U_{q}^{+}(L{\mathfrak{n}}^{-}):=U_{q}(L{\mathfrak{n}}^{-})\cap U_{q}(\widehat{{\mathfrak{b}}}^{+})=\Big{\{}\text{subalgebra generated by }{\mathsf{e}}_{\tilde{\beta}_{k}},k>0\Big{\}},$
(5.40) $U_{q}^{+}(L{\mathfrak{n}}^{+}):=U_{q}(L{\mathfrak{n}}^{+})\cap U_{q}(\widehat{{\mathfrak{b}}}^{+})=\Big{\{}\text{subalgebra generated by }{\mathsf{e}}_{\tilde{\beta}_{k}},k\leq 0\Big{\}},$
@@
(5.41) ${\mathsf{e}}_{\tilde{\beta}_{k}}=\begin{cases}\varphi_{-\text{hdeg }(\tilde{\beta}_{k})}E_{\tilde{\beta}_{k}}&\text{if }k>0\\ E_{\tilde{\beta}_{k}}&\text{if }k\leq 0\end{cases}\,.$
@@
$\varphi_{\pm\text{hdeg }(z)}:=\varphi_{\pm\sum_{i\in I}k_{i}\alpha_{i}}=\prod_{i\in I}\varphi^{\pm k_{i}}_{i}\,\in U_{q}(L{\mathfrak{h}})\,.$
@@
(5.46) ${\mathsf{e}}_{-\tilde{\beta}_{k}}=\Omega({\mathsf{e}}_{\tilde{\beta}_{k}})=\begin{cases}E_{-\tilde{\beta}_{k}}\varphi_{\text{hdeg }(\tilde{\beta}_{k})}&\text{if }k>0\\ E_{-\tilde{\beta}_{k}}&\text{if }k\leq 0\end{cases}\,.$
@@
where the sum is finite as it is taken over all tuples $n_{r-1},\dots,n_{s+1}\in{\mathbb{Z}}_{\geq 0}$ such that:
@@
A similar result also holds for $U_{q}(L{\mathfrak{n}}^{+})$ with ${\mathsf{e}}_{-\beta_{s}}$ replaced by ${\mathsf{e}}_{\beta_{s}}$.

### v2

@@ 4.3. Affine Weyl groups
$s_{i}(\mu)=\mu-(\alpha_{i},\mu)\alpha_{i}^{\vee}\ \qquad\text{and}\qquad s_{i}(\lambda)=\lambda-(\lambda,\alpha^{\vee}_{i})\alpha_{i}\,.$
@@
$s_{i}(\lambda,d)=\left(\lambda-(\lambda,\alpha^{\vee}_{i})\alpha_{i},d\right)\quad\forall\,i\in I\,,$
$s_{0}(\lambda,d)=\left(\lambda-(\lambda,\theta^{\vee})\theta,d+(\lambda,\theta^{\vee})\right)$
@@ 4.4. Reduced decompositions
Recall that the length of an element $x\in\widehat{W}$, denoted by $l(x)\in{\mathbb{N}}$, is the smallest number $l\in{\mathbb{N}}$ such that we can write $x=s_{i_{1-l}}\dots s_{i_{0}}$ for various $i_{1-l},\dots,i_{0}\in\widehat{I}$. Every such factorization is called a reduced decomposition of $x$. Given such a reduced decomposition, the terminal subset of the affine root system is:
@@ 4.6. Identification of orders
give convex orders of the sets $\Delta^{+}\times{\mathbb{Z}}_{<0}$ and $\Delta^{+}\times{\mathbb{Z}}_{\geq 0}$, respectively. We note that $\beta_{k+l}=\widehat{\mu}(\beta_{k})$ for any $k\in{\mathbb{Z}}$. Due to (4.9), if $\beta_{k}=(\alpha,d)$ and $\beta_{k+l}=(\alpha^{\prime},d^{\prime})$, then
(4.19) $\alpha=\alpha^{\prime}\quad\text{ and }\quad d=d^{\prime}+(\alpha,\mu)\,.$
This reveals a periodicity of the entire set $\Delta^{+}\times{\mathbb{Z}}$, not just $\Delta^{+}\times{\mathbb{Z}}_{<0}$ and $\Delta^{+}\times{\mathbb{Z}}_{\geq 0}$.
@@
For any $(\alpha,d)\in L_{N}$, the relative exponent of the first letter in $\ell(\alpha,-d)$ lies in $(-N-1;-N]$, due to Corollary 3.11 and Proposition 3.15. Thus, for any $(\alpha,d)\in L_{M}$, $(\beta,e)\in L_{N}$ with $M>N$, we have $\ell(\alpha,-d)>\ell(\beta,-e)$. As for the affine roots from the same block, consider $\beta_{r-Nl},\beta_{s-Nl}\in L_{N}$ with $1-l\leq s<r\leq 0$. If $\beta_{r}=(\alpha,d)$ and $\beta_{s}=(\beta,e)$, then $\beta_{r-Nl}=(\alpha,d+N\cdot f(\alpha))$ and $\beta_{s-Nl}=(\beta,e+N\cdot f(\beta))$, due to (4.19). On the other hand, the words $\ell(\alpha,-d-N\cdot f(\alpha))$ and $\ell(\beta,-e-N\cdot f(\beta))$ are obtained from $\ell(\alpha,-d)$ and $\ell(\beta,-e)$, respectively, by decreasing each relative exponent by $N$, due to Proposition 3.15. Since the latter operation obviously preserves the lexicographic order, and $\ell(\alpha,-d)<\ell(\beta,-e)$ as a consequence of $r>s$, we obtain the required inequality $\ell(\alpha,-d-N\cdot f(\alpha))<\ell(\beta,-e-N\cdot f(\beta))$. ∎
@@
$\dots<\beta_{3}<\beta_{2}<\beta_{1}<\beta_{0}<\beta_{-1}<\beta_{-2}<\cdots$
@@
$L_{<(i,d)}=\Big{\{}(\alpha,p)\,\Big{|}\,\alpha\in\Delta^{+},\,p\in{\mathbb{Z}}_{\geq 0},\,\ell(\alpha,-p)<\ell(\alpha_{i},-d)\Big{\}}.$
@@
First, let us prove that $\ell(\alpha,-p)<\ell(\alpha_{i},-d)=i^{(-d)}$ implies $p<p(\alpha)$. Let $j^{(-e)}$ be the first letter of $\ell(\alpha,-p)$, so that $j^{(-e)}<i^{(-d)}$. Hence, $e/c_{j}\leq d/c_{i}$ and the inequality is strict if $j\geq i$. This is equivalent to $e<p_{j}$, due to the definition (4.23). Then for any letter $\imath^{(-s)}\in\ell(\alpha,-p)$, we have $\imath^{(-s+1)}\leq j^{(-e)}<i^{(-d)}$ with the first inequality due to Theorem 3.6. As above, this implies $s-1<p_{\imath}$, so that $s\leq p_{\imath}$. Summing all these inequalities, we obtain the desired inequality $p<p(\alpha)$.
Let us prove the opposite implication by contradiction: assume that $\ell(\alpha,-p)>\ell(\alpha_{i},-d)$ for some $\alpha\in\Delta^{+}$ and $p<p(\alpha)$. Let $j^{(-e)}$ be the first letter of $\ell(\alpha,-p)$, so that $j^{(-e)}\geq i^{(-d)}$. As $\ell(\alpha,-p)$ is Lyndon, any letter $\imath^{(-s)}\in\ell(\alpha,-p)$ satisfies $j^{(-e)}\leq\imath^{(-s)}$. Therefore, $s/c_{\imath}\geq d/c_{i}$ and the inequality is strict for $\imath<i$. Thus, $s\geq p_{\imath}$. Summing all these inequalities, we obtain $p\geq p(\alpha)$, a contradiction. This completes our proof of $\ell(\alpha,-p)<\ell(\alpha_{i},-d)$ for any $0\leq p<p(\alpha)$. ∎
@@ 5.1. Quantum groups
If we write $\varphi_{i}=q_{i}^{h_{i}}$ and take the limit $q\rightarrow 1$, then $U_{q}({\mathfrak{g}})$ degenerates to $U({\mathfrak{g}})$. It is thus natural that many features of the latter also admit $q$-deformations. For example, let us recall the notion of standard Lyndon words from Subsections 2.12.7, and consider the following $q$-version of Definition 2.8 and the construction (2.4):
@@
$\begin{split}&U_{q}({\mathfrak{n}}^{+})\ =\bigoplus^{k\in{\mathbb{N}}}_{\ell_{1}\geq\dots\geq\ell_{k}\text{ standard Lyndon words}}{\mathbb{Q}}(q)\cdot e_{\ell_{1}}\dots e_{\ell_{k}}\ =\\ &\qquad\qquad\qquad\qquad\qquad\bigoplus_{w\text{--standard words}}{\mathbb{Q}}(q)\cdot e_{w}\,.\end{split}$
@@ 5.5. Quantum loop algebras
$\text{hdeg }x=\alpha\qquad\text{and}\qquad\text{vdeg }x=d\,,$
@@
$e_{w}=e_{\ell_{1}}\dots e_{\ell_{k}}\qquad\text{and}\qquad f_{w}=f_{\ell_{1}}\dots f_{\ell_{k}}$
@@
The analogous result also holds with $U_{q}(L{\mathfrak{n}}^{+})\leftrightarrow U_{q}(L{\mathfrak{n}}^{-})$ and $e\leftrightarrow f$.
@@ 5.9. Quantum affine algebras
We may define the ($q$-deformed) root vectors:
@@
$\Omega\colon e_{i}\mapsto f_{i},\ f_{i}\mapsto e_{i},\ \varphi^{\pm 1}_{i}\mapsto\varphi_{i}^{\mp 1},\ q\mapsto q^{-1}\qquad\forall\,i\in\widehat{I}\,.$
@@
[…]
@@ 5.13. Interplay of two algebras
It suffices to prove only (5.28) while (5.29) then follows as $\Omega$ commutes with the extended affine braid group action (due to (5.21) and $\Omega\circ T_{\tau}=T_{\tau}\circ\Omega$ for $\tau\in{\mathcal{T}}$).
@@
$T_{\widehat{\omega_{i,d}}}=\prod_{j\neq i}T^{p_{j}}_{\widehat{\omega_{j}^{\vee}}}\cdot T^{p_{i}}_{\widehat{\omega_{i}^{\vee}}}=\prod_{j\neq i}T^{p_{j}}_{\widehat{\omega_{j}^{\vee}}}\cdot T^{d}_{\widehat{\omega_{i}^{\vee}}}\,.$
@@ 5.15. PBW-type bases via quarter subalgebras
$U_{q}^{+}(L{\mathfrak{n}}^{-}):=U_{q}(L{\mathfrak{n}}^{-})\cap U_{q}(\widehat{{\mathfrak{b}}}^{+})=\Big{\{}\text{subalgebra generated by }{\mathsf{e}}_{\tilde{\beta}_{k}},k>0\Big{\}},$
$U_{q}^{+}(L{\mathfrak{n}}^{+}):=U_{q}(L{\mathfrak{n}}^{+})\cap U_{q}(\widehat{{\mathfrak{b}}}^{+})=\Big{\{}\text{subalgebra generated by }{\mathsf{e}}_{\tilde{\beta}_{k}},k\leq 0\Big{\}},$
@@
(5.32) ${\mathsf{e}}_{\tilde{\beta}_{k}}=\begin{cases}\varphi_{-\text{hdeg }\tilde{\beta}_{k}}E_{\tilde{\beta}_{k}}&\text{if }k>0\\ E_{\tilde{\beta}_{k}}&\text{if }k\leq 0\end{cases}\,.$
@@
$\varphi_{\pm\text{hdeg }z}:=\varphi_{\pm\sum_{i\in I}k_{i}\alpha_{i}}=\prod_{i\in I}\varphi^{\pm k_{i}}_{i}\,\in U_{q}(L{\mathfrak{h}})\,.$
@@
(5.37) ${\mathsf{e}}_{-\tilde{\beta}_{k}}=\Omega({\mathsf{e}}_{\tilde{\beta}_{k}})=\begin{cases}E_{-\tilde{\beta}_{k}}\varphi_{\text{hdeg }\tilde{\beta}_{k}}&\text{if }k>0\\ E_{-\tilde{\beta}_{k}}&\text{if }k\leq 0\end{cases}\,.$
@@
where the sum is finite as it is taken over all tuples $n_{r-1},\dots,n_{s+1}\in{\mathbb{N}}$ such that:
@@
The analogous result also holds for $U_{q}(L{\mathfrak{n}}^{+})$ with ${\mathsf{e}}_{-\beta_{s}}$ replaced by ${\mathsf{e}}_{\beta_{s}}$.
@@
6. Generalization to other orders
In this Section, we generalize our main results to a larger family of orders on the alphabet =${\mathcal{I}}=\{i^{(d)}\}_{i\in I}^{d\in{\mathbb{Z}}}$. Consider a collection of functions $\mathsf{f}_{i}\colon{\mathbb{Z}}\to{\mathbb{R}}$ such that
  • $\bullet$
    $\mathsf{f}_{i}(0)=0$
  • $\bullet$
    all $\mathsf{f}_{i}$ are strictly increasing unbounded functions
  • $\bullet$
    there are infinitely many $N$ (both in ${\mathbb{R}}_{>0}$ and ${\mathbb{R}}_{<0}$) such that there exist $\{N_{i}\}_{i\in I}\subset{\mathbb{Z}}^{I}$ satisfying $\mathsf{f}_{i}(N_{i})=N$ for all $i\in I$.
We then define an order on ${\mathcal{I}}$ (hence a lexicographic order on the loop words) via:
(6.1) $i^{(d)}<j^{(e)}\qquad\Longleftrightarrow\qquad\mathsf{f}_{i}(d)>\mathsf{f}_{j}(e)\quad\text{ or }\quad\mathsf{f}_{i}(d)=\mathsf{f}_{j}(e)\text{ and }i<j\,.$
In the special case $\mathsf{f}_{i}(d)=\frac{d}{c_{i}}$ (with $c_{i}\in{\mathbb{Z}}_{>0}$) this recovers (1.9) considered above.
$\bullet$ First, we need to update the filtration (3.4) of the loop algebra $L{\mathfrak{n}}^{+}$. To this end, we fix an increasing sequence $\{N^{(+,s)}\}_{s\in{\mathbb{N}}}$ of non-negative numbers (respectively, a decreasing sequence $\{N^{(-,s)}\}_{s\in{\mathbb{N}}}$ of non-positive numbers) such that $N^{(\pm,0)}=0$ and there exist $\{N^{(\pm,s)}_{i}\}_{i\in I}\subset(\pm{\mathbb{N}})^{I}$ satisfying $\mathsf{f}_{i}(N^{(\pm,s)}_{i})=N^{(\pm,s)}$ for all $i\in I$. Then, we define $L^{(s)}{\mathfrak{n}}^{+}$ as the finite-dimensional Lie subalgebra of $L{\mathfrak{n}}^{+}$ generated by
$\left\{e_{i}^{(d)}\,\Big{|}\,i\in I,\,N^{(-,s)}_{i}\leq d\leq N^{(+,s)}_{i}\right\}.$
We also amend our former definition of ${\mathcal{I}}^{(s)}$ in (3.3) by rather redefining
${\mathcal{I}}^{(s)}=\left\{i^{(d)}\,\Big{|}\,i\in I,\,N^{(-,s)}_{i}\leq d\leq N^{(+,s)}_{i}\right\}\qquad\forall\,s\in{\mathbb{N}}\,.$
We may thus apply Definition 2.10 to yield a notion of standard (Lyndon) loop words with respect to $L^{(s)}{\mathfrak{n}}^{+}$, with the words made up only of $i^{(d)}\in{\mathcal{I}}^{(s)}$.
$\bullet$ For $N$ as above, so that there exist $\{N_{i}\}_{i\in I}$ satisfying $N=\mathsf{f}_{i}(N_{i})\ \forall\,i$, we define
$f_{N}\colon\Delta^{+}\to{\mathbb{Z}}\quad\mathrm{via}\quad f_{N}(\alpha)=\sum_{i\in I}k_{i}\cdot N_{i}\quad\mathrm{for\ any}\quad\alpha=\sum_{i\in I}k_{i}\alpha_{i}\in\Delta^{+}\,.$
With this definition at hand, the subalgebra $L^{(s)}{\mathfrak{n}}^{+}$ can be explicitly written as
$L^{(s)}{\mathfrak{n}}^{+}=\bigoplus_{\alpha\in\Delta^{+}}\bigoplus_{d=f_{N^{(-,s)}}(\alpha)}^{f_{N^{(+,s)}}(\alpha)}{\mathbb{Q}}\cdot e_{\alpha}^{(d)}\,.$
Then, the results of Proposition 3.2 and Proposition 3.3 still hold true with the only change that $-sf(\alpha)\leq d\leq sf(\alpha)$ is replaced with $f_{N^{(-,s)}}(\alpha)\leq d\leq f_{N^{(+,s)}}(\alpha)$.
$\bullet$ As before, we call a loop word $w=\left[i_{1}^{(d_{1})}\dots\,i_{n}^{(d_{n})}\right]$ exponent-tight if (3.9) holds. Then, the results of Theorem 3.6, Lemma 3.7, Proposition 3.8, and Proposition 3.13 still hold true (the proofs are the same). Therefore, we still have a bijection (3.13)
[…]
$\bullet$ On the other hand, the periodicity of Proposition 3.15 no longer holds in this generality. Instead, we can only express $\ell(\alpha,f_{N}(\alpha))$ via $\ell(\alpha,0)$:
Lemma 6.1.
If $\ell(\alpha,0)=[i_{1}^{(0)}\dots\,i_{n}^{(0)}]$, then $\ell(\alpha,f_{N}(\alpha))=[i_{1}^{(N_{i_{1}})}\dots\,i_{n}^{(N_{i_{n}})}]$.
Proof.
First, we note that Theorem 3.6 together with Proposition 3.8 for $N>0$ (respectively, Remark 3.9 for $N<0$) guarantee that the multiset of letters constituting $\ell(\alpha,f_{N}(\alpha))$ is exactly $\{i_{1}^{(N_{i_{1}})},\dots,i_{n}^{(N_{i_{n}})}\}$. Indeed, assuming the contradiction for some $N>0$ (the case $N<0$ is treated analogously), there exists $0\leq d<f_{N}(\alpha)$ such that $\ell(\alpha,d+1)$ starts with $i_{k}^{(N_{i_{k}}+1)}$ for some $k$. As the sum of exponents equals $d+1\leq f_{N}(\alpha)$, the word $\ell(\alpha,d+1)$ and hence $\ell(\alpha,d)$ also contains a letter $i_{l}^{(e)}$ with $e<N_{i_{l}}$. This provides a contradiction with Proposition 3.8, since $i_{l}^{(e+1)}>i_{k}^{(N_{i_{k}}+1)}$.
On the other hand, we note that $i^{(N_{i})}<j^{(N_{j})}$ iff $i<j$, which guarantees that the loop word $[j_{1}^{(N_{j_{1}})}\dots\,j_{n}^{(N_{j_{n}})}]$ is (standard) Lyndon iff the loop word $[j_{1}^{(0)}\dots\,j_{n}^{(0)}]$ is (standard) Lyndon, cf. the proof of Proposition 3.15. This completes the proof. ∎
We can now prove the following slight generalization of Corollary 3.11:
Lemma 6.2.
Fix $N$, $\{N_{i}\}_{i\in I}$, $\alpha\in\Delta^{+}$, $f_{N}(\alpha)\in{\mathbb{Z}}$ as above.
(a) For $d>f_{N}(\alpha)$, the loop word $\ell(\alpha,d)$ starts with some $j^{(e)}$ such that $e>N_{j}$.
(b) For $d\leq f_{N}(\alpha)$, the loop word $\ell(\alpha,d)$ starts with some $j^{(e)}$ such that $e\leq N_{j}$.
Proof.
Let $\ell(\alpha,d)=[i_{1}^{(d_{1})}\dots\,i_{n}^{(d_{n})}]$. Then $i_{1}^{(d_{1})}\leq i_{r}^{(d_{r})}$ and so $\mathsf{f}_{i_{1}}(d_{1})\geq\mathsf{f}_{i_{r}}(d_{r})$ for any $r$. If $d_{1}\leq N_{i_{1}}$, then we would have $\mathsf{f}_{i_{r}}(N_{i_{r}})=\mathsf{f}_{i_{1}}(N_{i_{1}})\geq\mathsf{f}_{i_{1}}(d_{1})\geq\mathsf{f}_{i_{r}}(d_{r})$ and so $d_{r}\leq N_{i_{r}}$ for any $r$. That would imply $d=\sum_{r=1}^{n}d_{r}\leq f_{N}(\alpha)$, a contradiction.
To prove (b), we note that $i_{1}^{(d_{1})}\geq i_{r}^{(d_{r}+1)}$ for any $r$ by Theorem 3.6. If $d_{1}>N_{i_{1}}$, then we would have $\mathsf{f}_{i_{r}}(N_{i_{r}})=\mathsf{f}_{i_{1}}(N_{i_{1}})<\mathsf{f}_{i_{1}}(d_{1})\leq\mathsf{f}_{i_{r}}(d_{r}+1)$ so that $d_{r}\geq N_{i_{r}}$ for all $r$. That would imply $d=\sum_{r=1}^{n}d_{r}>f_{N}(\alpha)$, a contradiction. ∎
$\bullet$ The major difference will take place in the generalization of Theorem 4.7 to the present setup. As the periodicity of Proposition 3.15 no longer holds, the bi-infinite sequence $\{i_{k}\}_{k\in{\mathbb{Z}}}$ of (4.15) shall rather be constructed as a limit of finite sequences. Explicitly, to define $\{i_{k}\}_{k\leq 0}$, instead of $L$ from (3.14) we shall consider
$L^{[s]}=\Big{\{}(\alpha,d)\,\Big{|}\,\alpha\in\Delta^{+},0\leq d<f_{N^{(+,s)}}(\alpha)\Big{\}}\qquad\forall\,s\in{\mathbb{Z}}_{>0}.$
These “blocks” can be identified with the following terminal sets:
$L^{[s]}=E_{\widehat{\mu^{(+,s)}}}\quad\mathrm{with}\quad\mu^{(+,s)}=\sum_{i\in I}N_{i}^{(+,s)}\omega^{\vee}_{i}.$
Arguing as in the proof of Theorem 4.7, we get a reduced decomposition (4.14) of $\widehat{\mu^{(+,s)}}$ such that the ordered finite sequence $\beta_{0}<\beta_{-1}<\dots<\beta_{1-l(\widehat{\mu^{(+,s)})}}$ coincides with $L^{[s]}$ ordered via (4.21). By uniqueness, such a sequence for $L^{[s+1]}$ refines the one for $L^{[s]}$. Furthermore, the roots $\beta_{k}=s_{i_{0}}s_{i_{-1}}\dots s_{i_{k+1}}(\alpha_{i_{k}})$ for $k\leq 0$ are all distinct and satisfy $\{\beta_{k}\}_{k\leq 0}=\Delta^{+}\times{\mathbb{Z}}_{\geq 0}$. The construction of $\{i_{k}\}_{k>0}$ is similar.
$\bullet$ With the above update of Theorem 4.7, the results of Propositions 4.94.10 still hold (for any fixed $i\in I,d\in{\mathbb{Z}}$), once $\{p_{j}\}_{j\in I}$ from (4.23) are rather redefined via:
$p_{j}\ \ \mathrm{is\ the\ unique\ integer\ satisfying}\ \ j^{(-p_{j})}\geq i^{(-d)}>j^{(-p_{j}+1)}\,$
with $i\in I,d\geq 0$ fixed. This recovers (4.23) if $\mathsf{f}_{j}(d)=\frac{d}{c_{j}}$ for all $j$ (with $c_{j}\in{\mathbb{Z}}_{>0}$).
$\bullet$ With the above update, the analogue of (5.31) and the paragraph afterwards hold, implying (5.285.29). Thus, the main result of Section 5, the construction of PBW bases of $U_{q}(L{\mathfrak{n}}^{+})$ from Theorem 5.8 still holds (with $e_{\ell},e_{w}$ as in Definition 5.7).